Stomach acid is approximately 0.137 M HCl. Using your average number of moles of CaCO3 per antacid tablet, what volume of stomach acid will a tablet neutralize?

1
Expert's answer
2022-04-07T04:08:04-0400


2 mol HCl - 1 mol CaCO3

x mol HCl - 4.955×10-3 mol CaCO3

x = 4.955×10-3 × 2 = 0.00991 mol

V(sol) = n(HCl) / C(sol) = 0.00991 mol / 0.137 mol/l = 0.07233 l = 72.33 ml